Grade 12Inorganic Chemistry

find normality of the following

1. 70 g of NH4OH is dissolved in 2 L of water

2 . 39 g of Al(OH)3 is dissolved in 10 L of water.

ANS) 1.) 1.12

2.)0.15

plz shw the steps....

Mohammed Patel
13 Years agoGrade 12

2 Answers

Prajwal kr
ApprovedApproved Tutor Answer13 Years ago

1. n factor = 1

    no. of moles of NH4OH is: 70/35=2.

    Molarity= 2/2=1 M

    => Normality=1*1=1 N.......(answer)

2. n factor=3

    no. of moles of aluminium hydroxide=2

    Molarity=0.2 M

    Normality=0.6N........(answer)

 

Akash Kumar Dutta
ApprovedApproved Tutor Answer13 Years ago

1)norm.=70/32.2=1.12..since n-factor=1

2)norm.39.3/79.10=.15..since n-factor=3